# Computational Modelling of Thixotropic Multiphase Fluids

### v1

@@ Abstract
[label1]organization=Escuela de Ingeniería de Bilbao, Universidad del País Vasco (EHU),addressline=Plaza Ingeniero Torres Quevedo, 1, city=Bilbao, postcode=48013, state=Basque Country, country=Spain
@@
[label2]organization=Basque Center for Applied Mathematics (BCAM),addressline=Alameda de Mazarredo, 14, city=Bilbao, postcode=48400, state=Basque Country, country=Spain
@@ 1 Introduction
Multiphase systems are crucial to engineering, biology, and materials science, where the existence of complex microstructural dynamics often lead to non-trivial rheological behavior, such as shear-thinning or thixotropic properties. In industrial applications, understanding emulsion and foam stability is critical for improving product quality and efficiency [63]. In biology, liquid-liquid phase separation within cells organizes cellular components and influences diseases like neurodegenerative disorders [19, 10, 21, 14]. Similarly, interactions among blood components, such as plasma and red blood cells, are essential for processes like clotting under varying physiological conditions [4, 9, 2]. In materials science, understanding phase interactions is crucial for designing smart materials, composites, and alloys with desired properties [64, 65]. Accurately modeling these systems, including phase separation dynamics and time-dependent material properties [62], remains challenging, driving the development of advanced grid-based and mesh-free methods.
@@
Grid-based methods such as the Finite Element [77, 78, 79], Lattice Boltzmann Method (LBM) [80, 81, 82], methods based on the Volume of Fluid (VOF) [72, 73, 71], Level Set (LS) [74, 75, 76] or Immersed Boundary Method (IBM) [97, 98], are widely used for tracking interfaces and capturing phase interactions. These methods are particularly effective for problems involving sharp interfaces. However, they can struggle with dynamic interface tracking, requiring significant computational resources for high-resolution simulations. Additionally, spurious non-physical velocity fields near the interfaces (resulting from discretization errors) can change the shape or stability of the simulated phases [66, 67, 68]. Mesh-free methods, such as Smoothed Particle Hydrodynamics (SPH) [25, 94, 95, 6] and Dissipative Particle Dynamics (DPD) [83, 96, 84, 85], have gained popularity for their ability to handle complex geometries and dynamic interfaces without the need for explicitly tracking the interfaces. These methods are particularly advantageous for simulating multiphase flows with large deformations[6] and complex interactions[87]. However, they can suffer from numerical instabilities, especially at interfaces, where large differences in viscosity and density can generate pressure oscillations and unphysical artifacts, such as gaps or particle mixing at the interface [8, 3, 7, 70, 5]. Additionally, implementing accurate boundary conditions at walls and interfaces for fluids with varying viscosities is complex and can result in errors affecting flow behavior near these regions.
@@
Smoothed dissipative particle dynamics (SDPD) [12, 11] is a mesoscopic technique that combines the discretization scheme of SPH and the consistent thermal fluctuations of DPD, effectively reproducing the fluctuating Navier-Stokes equations. SDPD has been successfully applied to model a range of synthetic [60, 20, 37] and biological systems [18, 26, 27]. Recently, Lei et al. [17] developed a SDPD scheme suitable to model multiphase systems incorporating surface tension effects via a pairwise interaction forces. This scheme was able to capture complex phenomena such as droplet coalescence and interface stabilization in the presence of thermal noise, showing results consistent with theoretical predictions and experimental data. The method’s ability to incorporate thermal fluctuations makes it particularly suitable for studying micro to nanoscale systems where thermal fluctuations can play a role in phase interactions and dynamics. However, the application of SDPD to multiphase flows with complex rheological behavior, such as thixotropic materials, which is crucial for accurately simulating complex fluids like gels or emulsions that exhibit microstructural changes during flow, remains an emerging field of research.
@@ 2 Numerical methodology
In SDPD, a fluid is discretized using particles with a volume $V_{i}$, leading to a particle number density $d_{i}=1/V_{i}=\sum_{j}W(r_{ij},h)$. Where $W(r_{ij},h)$ is a kernel function that depends on the distance $r_{ij}=|{\bf{r}}_{i}-{\bf{r}}_{j}|$ between particles $i$ and $j$, has a finite support $h$ and is normalized to one. The evolution of the particle position is given by the equation ${\text{d}{\bf{r}}_{i}}/{\text{d}t}={\bf v}_{i}$, where ${\bf v}_{i}$ is the velocity of the $i$-th particle. The momentum’s stochastic differential equation is given by
$m_{i}\frac{\text{d}{\bf v}_{i}}{\text{d}t}$ $=-\sum_{j}\left[\frac{p_{i}}{d_{i}^{2}}+\frac{p_{j}}{d_{j}^{2}}\right]F_{ij}{\bf r}_{ij}$ (1)
$-\sum_{j}\left[a_{ij}{\bf v}_{ij}+b_{ij}({\bf v}_{ij}\cdot{{\bf e}_{ij}}){{\bf e}_{ij}}\right]\frac{F_{ij}}{d_{i}d_{j}}$
@@
where $p_{i}$ and $p_{j}$ are the pressure of particles $i$ and $j$, respectively. The pressure of each particle is determined by a suitable density-dependent equation of state $p=f(\rho)$, as discussed later. The term $m_{i}=\rho\Delta^{D}$, denotes the mass of the particle that depends on the density $\rho$, the equilibrium interparticle distance $\Delta$ and the dimension of the system $D$. The term $F_{ij}=-\nabla W(r_{ij},h)/r_{ij}$ is a positive function introduced to account for the interaction forces between particles. Here, we adopt a Lucy’s Quartic kernel [28] of the form $W(r)=\mathrm{w_{0}}(1+3\frac{r}{h})(1-\frac{r}{h})^{3}$ if $r/h\leq 1$ and $W(r)=0$ if $r/h>1$, where $\mathrm{w_{0}}=5/(\pi h^{2})$ for two dimensions and $\mathrm{w_{0}}=105/(16\pi h^{3})$ for three dimensions. The second term in (2) corresponds to viscous contributions, where ${\bf v}_{ij}={\bf v}_{i}-{\bf v}_{j}$ is the relative velocity between particles, ${{\bf e}_{ij}}={\bf r}_{ij}/|{\bf r}_{ij}|$ is the unit vector, $a_{ij}$ and $b_{ij}$ are friction coefficients related to the pairs shear viscosity ($\eta_{ij}$) and bulk viscosity ($\zeta_{ij}=\eta_{ij}(2D-1)/D$), with $a_{ij}={(D+2)\eta_{ij}}/{D}-\zeta_{ij}$ and $b_{ij}=(D+2)(\zeta_{ij}+{\eta_{ij}}/{D})$. Here, we consider the viscosity between particles $i$ and $j$ expressed as $\eta_{ij}=2(\eta_{i}\eta_{j})/(\eta_{i}+\eta_{j})$ The last term in (2) incorporates thermal fluctuations into the momentum balance by including a matrix of independent increments of the Wiener process, $d{\bf{W}}_{ij}$ and its traceless symmetric part $d{\bf\overline{W}}_{ij}$ given by
@@
The parameters $A_{ij}$ and $B_{ij}$ are the amplitude of the noise, related to the friction coefficients ($a_{ij}$ and $b_{ij}$) and is given by
$A_{ij}$ $=\left[4\ k_{B}T\ a_{ij}\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2}$ (4)
$B_{ij}$ $=\left[4\ k_{B}T\ \left(b_{ij}-a_{ij}\frac{D-2}{D}\right)\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2},$ (5)
@@ 2.1 Thixotropic model
When exposed to continuous shear stress, some non-Newtonian fluids can exhibit a time-dependent shear-thinning or thixotropic behavior, which is characterized by a gradual decrease in viscosity followed by a recovery upon removal of the stress. To capture this behavior, we can follow the phenomenological description given by Le-Cao et al. [16], where viscosity of a fluid is given by $\eta(t,x)=\eta_{\infty}\left[1+\alpha f(t,\dot{\gamma}(x))\right]$, where $\eta(t)$ is the time-dependent shear viscosity, $\eta_{\infty}$ is the limiting viscosity, $\alpha$ is a constitutive constant that determines the maximum viscosity and $f=[0:1]$, as indicated by Rossi et al. [62], is a microstructural scalar parameter describing the current state of the microstructure from fully destroyed ($f=0$) to completely developed ($f=1$). The time evolution of $f$ is given by $\dot{f}=a-(a+b\dot{\gamma})f$, where $a$ is a constant that represents the rate of formation of the microstructure, and $b\dot{\gamma}$ the rate of destruction of the microstructure under shear conditions. The strain-rate tensor $\dot{\boldsymbol{\gamma}}$ is given by $\dot{\boldsymbol{\gamma}}=(\nabla{\bf v})+(\nabla{\bf v}^{T})$. The strain-rate tensor second invariant $\dot{\gamma}$ can be evaluated as:
@@
The velocity gradient tensor $\nabla{\bf v}$ is given by
$(\nabla{\bf v})^{\alpha\beta}_{i}=\sum_{j}\frac{({\bf v}^{\alpha}_{i}-{\bf v}^{\alpha}_{j})({\bf r}_{i}^{\beta}-{\bf r}_{j}^{\beta})}{d_{j}r_{ij}}\frac{\partial W(r_{ij},h)}{\partial r},$ (7)
where $\alpha,\beta$ are the column and row indices in the matrix tensor and $\partial W(r_{ij},h)/\partial r$ the gradient of the kernel. By considering the ratio between the formation and destruction and the characteristic time for microstructure formation, we can introduce the parameter $\beta=b/a$ and the thixotropic time scale $\lambda_{0}=1/a$. Accordingly, the temporal evolution can be rewritten as $\dot{f}=[1-(1+\beta\dot{\gamma})f]\lambda_{0}^{-1}$ and, after time integration, the scalar parameter $f$ can be calculated as
$f(t)=\frac{1}{1+\beta\dot{\gamma}}\left(1-\exp^{-\frac{(1+\beta\dot{\gamma})t}{\lambda_{0}}}\right)+f_{0}\exp^{-\frac{(1+\beta\dot{\gamma})t}{\lambda_{0}}},$ (8)
where $f_{0}$ is the initial value of the scalar parameter $f$. Similar to the viscosity between a pair of particles $i$ and $j$, any pair parameter $X_{ij}$ of the thixotropic model is approximated by $X_{ij}=2(X_{i}X_{j})/(X_{i}+X_{j})$. For instance, we have for the limiting viscosity $(\eta_{\infty})_{ij}=2((\eta_{\infty})_{i}(\eta_{\infty})_{j})/((\eta_{\infty})_{i}+(\eta_{\infty})_{j})$, and similarly for the rest of factors.
@@ 2.2 Multiphase Model
Interfacial tension can be included in the model using an additional pairwise interaction force $\mathbf{F}^{int}_{ij}$, to the momentum balance (2). This interaction force takes the form[25]
$\mathbf{F}^{int}_{ij}=\mathbf{F}^{int}(\mathbf{r}_{ij})=-s_{ij}\phi(r_{ij})(\mathbf{r}_{ij}/r_{ij}),$ (9)
where $s_{ij}$ and $\phi$ determine the strength and functional form of the potential, respectively. This force is short-range repulsive and long-range attractive [17]. For two different phases $\alpha$ and $\beta$ the factor $s_{ij}$ is related to the domains $\Omega_{\alpha}$ and $\Omega_{\beta}$ as
$s_{ij}=\begin{cases}s_{\alpha\alpha}&\text{if }\mathbf{r}_{i}\in\Omega_{\alpha}\text{ and }\mathbf{r}_{j}\in\Omega_{\alpha},\\ s_{\beta\beta}&\text{if }\mathbf{r}_{i}\in\Omega_{\beta}\text{ and }\mathbf{r}_{j}\in\Omega_{\beta},\\ s_{\alpha\beta}&\text{if }\mathbf{r}_{i}\in\Omega_{\alpha}\text{ and }\mathbf{r}_{j}\in\Omega_{\beta},\end{cases}$ (10)
@@
$s_{\alpha\alpha}=s_{\beta\beta}=10^{k}s_{\alpha\beta}=\frac{1}{1-10^{-k}}n^{-2}_{eq}\frac{\sigma_{0}}{\lambda_{m}}$ (11)
Here, $k$ is a constant that must be greater than 1, $\sigma_{0}$ is a macroscopic surface tension. $n_{eq}=\Delta^{-D}$ is the equilibrium particle density which depends on the grid size $\Delta$. The factors $\lambda_{m}$ and $\phi(r_{ij})$ are given by
$\lambda_{m}$ $=\sqrt{2\pi}(-A\epsilon_{0}^{5}+\epsilon^{5}),$ (12)
$\phi(r_{ij})$ $=r_{ij}\left[-Ae^{-\frac{r^{2}_{ij}}{2\epsilon_{0}^{2}}}+e^{-\frac{r^{2}_{ij}}{2\epsilon^{2}}}\right],$ (13)
where $A$ represents the magnitude of the ratio between repulsive and attractive forces and $\epsilon$ and $\epsilon_{0}$ are compact-support dependent functions. Following Tartakovsky and Panchenko [25], we use $\epsilon=h/3.5$, $\epsilon_{0}=\epsilon/2$ and $A=(\epsilon/\epsilon_{0})^{3}$. This choice ensures that the virial pressure ($P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})n_{eq}^{2}s_{\alpha\alpha}$) is negative, which is necessary for the stability of the interfaces [25]. However, we must note, that the presence of a negative virial pressure can also lead to numerical instabilities in the system, especially when the system is not in equilibrium. In order to prevent this viral pressure affects the global system pressure, we introduce an equation of state fo the form
@@
where $\rho_{0}$ is the equilibrium density for the system, $c_{s}$ is the speed of sound, and $P_{b}=-P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})n_{eq}^{2}s_{\alpha\alpha}$ is a background pressure that ensures $P>0$. This component balance the system and avoid spurious effects of global negative pressure. It is important to note that the presence of this background pressure can also influence the dynamics of the system, especially in flow conditions. In the Results section, we show that the choice of $P_{b}$ does not introduce artifacts into the simulations, under different dynamic conditions.
In our simulations, we adopt the well-known Velvet-velocity algorithm [36] for the temporal evolution of the position and velocity of the particles. We use a time step $\Delta t$ that ensure numerical stability[37], satisfying both: $\Delta t\leq 0.125h^{2}\rho/\eta$ (the viscous diffusion criteria [38]) and $\Delta t\leq 0.25h/c_{s}$ (the Courant-Friedrichs-Lewy (CFL) condition [39]).

### v2

@@ Abstract
[label1]organization=Basque Center for Applied Mathematics (BCAM),addressline=Alameda de Mazarredo, 14, city=Bilbao, postcode=48400, state=Basque Country, country=Spain
@@
[label2]organization=University of the Basque Country/Euskal Herriko Unibertsitatea,addressline=Barrio Sarriena, city=Leioa, postcode=48940, state=Basque Country, country=Spain
@@ 1 Introduction
Multiphase systems are crucial to engineering, biology, and materials science, where the existence of complex microstructural dynamics often lead to non-trivial rheological behavior, such as shear-thinning or thixotropic properties. In industrial applications, understanding emulsion and foam stability is critical for improving product quality and efficiency dickinson2020advances . In biology, liquid-liquid phase separation within cells organizes cellular components and influences diseases like neurodegenerative disorders shin2017liquid , alberti2021biomolecular , spannl2019biomolecular , hyman2014liquid . Similarly, interactions among blood components, such as plasma and red blood cells, are essential for processes like clotting under varying physiological conditions fogelson2015fluid , smith2015all , ataullakhanov1998spatiotemporal . In materials science, understanding phase interactions is crucial for designing smart materials, composites, and alloys with desired properties huang2021interphase , pramod2015aluminum . Accurately modeling these systems, including phase separation dynamics and time-dependent material properties rossi2022sph , remains challenging, driving the development of advanced grid-based and mesh-free methods.
@@
Grid-based methods such as the Finite Element soloveichik2022method , molina2023finite , mavridis1987finite , Lattice Boltzmann Method (LBM) yan2011lbm , petersen2021lattice , reis2022lattice , methods based on the Volume of Fluid (VOF) dolai2025hybrid , garoosi2022numerical , wang2022volume , Level Set (LS) jettestuen2021locally , doherty2023stabilised , lyras2023conservative or Immersed Boundary Method (IBM) patel2017coupled , zhenga2017coupled , are widely used for tracking interfaces and capturing phase interactions. These methods are particularly effective for problems involving sharp interfaces. However, they can struggle with dynamic interface tracking, requiring significant computational resources for high-resolution simulations. Additionally, spurious non-physical velocity fields near the interfaces (resulting from discretization errors) can change the shape or stability of the simulated phases leclaire2015approach , xu2017lattice , leclaire2016modeling . Mesh-free methods, such as Smoothed Particle Hydrodynamics (SPH) Ellero2007 , tartakovsky2016pairwise , hu2009constant , le2025smoothed , liu2010smoothed and Dissipative Particle Dynamics (DPD) GrootAndWarren1997 , espanol2017perspective , gavrilov2011microphase , tiwari2006dissipative , have gained popularity for their ability to handle complex geometries and dynamic interfaces without the need for explicitly tracking the interfaces. These methods are particularly advantageous for simulating multiphase flows with large deformationsliu2010smoothed and complex interactionsmoreno2020 . However, they can suffer from numerical instabilities, especially at interfaces, where large differences in viscosity and density can generate pressure oscillations and unphysical artifacts, such as gaps or particle mixing at the interface monaghan2013simple , chen2015sph , meng2020multiphase , yang2022consistent , guo2024smoothed . Additionally, implementing accurate boundary conditions at walls and interfaces for fluids with varying viscosities is complex and can result in errors affecting flow behavior near these regions.
@@
Smoothed dissipative particle dynamics (SDPD) espanol2003smoothed , ellero2018everything is a mesoscopic technique that combines the discretization scheme of SPH and the consistent thermal fluctuations of DPD, effectively reproducing the fluctuating Navier-Stokes equations. SDPD has been successfully applied to model a range of synthetic vazquez2012sph , simavilla2022mesoscopic , bian2012multiscale , moreno2023 , echeverria2025 and biological systems moreno2013multiscale , zohravi2023computational , zohravi2025mesoscale , moreno-chaparro2025 . Recently, Lei et al. lei2016smoothed developed a SDPD scheme suitable to model multiphase systems incorporating surface tension effects via a pairwise interaction forces. This scheme was able to capture complex phenomena such as droplet coalescence and interface stabilization in the presence of thermal noise, showing results consistent with theoretical predictions and experimental data. The method’s ability to incorporate thermal fluctuations makes it particularly suitable for studying micro to nanoscale systems where thermal fluctuations can play a role in phase interactions and dynamics. However, the application of SDPD to multiphase flows with complex rheological behavior, such as thixotropic materials, which is crucial for accurately simulating complex fluids like gels or emulsions that exhibit microstructural changes during flow, remains an emerging field of research.
@@ 2 Numerical methodology
In SDPD, a fluid is discretized using particles with a volume $V_{i}$, leading to a particle number density $d_{i}=1/V_{i}=\sum_{j}W(r_{ij},h)$. $W(r_{ij},h)$ is a kernel function that depends on the distance $r_{ij}=|{\bf{r}}_{i}-{\bf{r}}_{j}|$ between particles $i$ and $j$, has a finite support $h$ and is normalized to one. The evolution of the particle position is given by the equation ${\text{d}{\bf{r}}_{i}}/{\text{d}t}={\bf v}_{i}$, where ${\bf v}_{i}$ is the velocity of the $i$-th particle. The momentum’s stochastic differential equation is given by
$m_{i}\frac{\text{d}{\bf v}_{i}}{\text{d}t}$ $=\sum_{j}\left[\frac{P_{i}}{d_{i}^{2}}+\frac{P_{j}}{d_{j}^{2}}\right]F_{ij}{\bf r}_{ij}$ (1)
$-\sum_{j}\left[\mathcal{A}_{ij}{\bf v}_{ij}+\mathcal{B}_{ij}({\bf v}_{ij}\cdot{{\bf e}_{ij}}){{\bf e}_{ij}}\right]\frac{F_{ij}}{d_{i}d_{j}}$
@@
where $P_{i}$ and $P_{j}$ are the pressure of particles $i$ and $j$, respectively. The pressure of each particle is determined by a suitable density-dependent equation of state $P=f(\rho)$, as discussed later. The term $m_{i}=\rho{\Delta x}^{D}$, denotes the mass of the particle that depends on the density $\rho$, the equilibrium interparticle distance $\Delta x$, and the dimension of the system $D$. The term $F_{ij}=-\nabla W(r_{ij},h)/r_{ij}$ is a positive function introduced to account for the interaction forces between particles. Here, we adopt a Lucy’s Quartic kernel lucy1977numerical of the form $W(r)=\mathrm{w_{0}}(1+3{r}/{h})(1-{r}/{h})^{3}$ if $r/h\leq 1$ and $W(r)=0$ if $r/h>1$, where $\mathrm{w_{0}}=5/(\pi h^{2})$ for two dimensions and $\mathrm{w_{0}}=105/(16\pi h^{3})$ for three dimensions. The second term in (2) corresponds to viscous contributions, where ${\bf v}_{ij}={\bf v}_{i}-{\bf v}_{j}$ is the relative velocity between particles, ${{\bf e}_{ij}}={\bf r}_{ij}/|{\bf r}_{ij}|$ is the unit vector, $\mathcal{A}_{ij}$ and $\mathcal{B}_{ij}$ are friction coefficients related to the pairs shear viscosity ($\eta_{ij}$) and bulk viscosity ($\zeta_{ij}=\eta_{ij}(2D-1)/D$), with $\mathcal{A}_{ij}={(D+2)\eta_{ij}}/{D}-\zeta_{ij}$ and $\mathcal{B}_{ij}=(D+2)(\zeta_{ij}+{\eta_{ij}}/{D})$. Here, we consider the viscosity between particles $i$ and $j$ expressed as $\eta_{ij}=2(\eta_{i}\eta_{j})/(\eta_{i}+\eta_{j})$. The last term in (2) incorporates thermal fluctuations into the momentum balance by including a matrix of independent increments of the Wiener process, $d{\bf{W}}_{ij}$ and its traceless symmetric part $d{\bf\overline{W}}_{ij}$ given by
@@
The parameters $A_{ij}$ and $B_{ij}$ are the amplitude of the noise, related to the friction coefficients ($\mathcal{A}_{ij}$ and $\mathcal{B}_{ij}$) and, in order to satisfy the fluctuation-dissipation theorem, they are given by
$A_{ij}$ $=\left[4\ k_{B}T\ \mathcal{A}_{ij}\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2}$ (4)
$B_{ij}$ $=\left[4\ k_{B}T\ \left(\mathcal{B}_{ij}-\mathcal{A}_{ij}\frac{D-2}{D}\right)\ \frac{F_{ij}}{d_{i}d_{j}}\right]^{-1/2},$ (5)
@@ 2.1 Thixotropic model
When exposed to continuous shear stress, some non-Newtonian fluids can exhibit a time-dependent shear-thinning or thixotropic behavior, which is characterized by a gradual decrease in viscosity followed by a recovery upon removal of the stress. To capture this behavior, we can follow the phenomenological description given by Le-Cao et al. le2020microstructure , where viscosity of a fluid is given by
$\eta(t,x)=\eta_{\infty}\left[1+\alpha f(t,\dot{\gamma}(x))\right]$ (6)
where $\eta(t)$ is the time-dependent shear viscosity, $\eta_{\infty}$ is the limiting lowest viscosity, $\alpha$ is a constitutive constant, and $f\in[0:1]$ is a microstructural scalar parameter describing the current state of the microstructure, continuously from fully destroyed ($f=0$) to completely developed ($f=1$) rossi2022sph . Notice from Equation 6, that for a fully developed microstructure, $\alpha$ indicates the maximum viscosity $\eta_{\text{max}}=\eta_{\infty}(1+\alpha)$. The time evolution of $f$ is given by
$\dot{f}=k_{f}-(k_{f}+k_{d}\dot{\gamma})f$ (7)
where $k_{f}$ is a constant that represents the rate of formation of the microstructure, and $k_{d}\dot{\gamma}$ the rate of destruction of the microstructure under shear conditions. The strain-rate tensor $\dot{\boldsymbol{\gamma}}$ is given by $\dot{\boldsymbol{\gamma}}=(\nabla{\bf v})+(\nabla{\bf v}^{T})$. The strain-rate tensor second invariant $\dot{\gamma}$ appearing in Equation 6 can be evaluated as:
@@
The velocity gradient tensor $\nabla{\bf v}$ is calculated in SDPD as
$(\nabla{\bf v})^{\alpha^{\prime}\beta^{\prime}}_{i}=\sum_{j}\frac{({\bf v}^{\alpha^{\prime}}_{i}-{\bf v}^{\alpha^{\prime}}_{j})({\bf r}_{i}^{\beta^{\prime}}-{\bf r}_{j}^{\beta^{\prime}})}{d_{j}r_{ij}}\frac{\partial W(r_{ij},h)}{\partial r},$ (9)
where $\alpha^{\prime},\beta^{\prime}$ are the column and row indices in the matrix tensor and $\partial W(r_{ij},h)/\partial r$ is the gradient of the kernel. By considering the ratio between the formation and destruction and the characteristic time for microstructure formation, we can introduce the parameter $\beta=k_{d}/k_{f}$ and the thixotropic time scale $\lambda_{\text{thix}}=1/k_{f}$. Accordingly, the temporal evolution can be rewritten as $\dot{f}=[1-(1+\beta\dot{\gamma})f]\lambda_{\text{thix}}^{-1}$ and, after time integration, the scalar parameter $f$ can be calculated as rossi2022sph
$f(t)=\frac{1}{1+\beta\dot{\gamma}}\left(1-\exp^{-{(1+\beta\dot{\gamma})t}/{\lambda_{\text{thix}}}}\right)+f_{0}\exp^{-{(1+\beta\dot{\gamma})t}/{\lambda_{\text{thix}}}},$ (10)
where $f_{0}$ is the initial value of the scalar parameter $f$. Similar to the viscosity between a pair of particles $i$ and $j$, any pair parameter $X_{ij}$ of the thixotropic model is approximated by $X_{ij}=2(X_{i}X_{j})/(X_{i}+X_{j})$. For instance, we have for the limiting viscosity $(\eta_{\infty})_{ij}=2((\eta_{\infty})_{i}(\eta_{\infty})_{j})/((\eta_{\infty})_{i}+(\eta_{\infty})_{j})$, and similarly for the rest of factors. A similar microstructural SPH approach has been recently considered to model also discontinuous shear thickening in flow of dense suspensions angerman2024microstructural , angerman2025numerical .
@@ 2.2 Multiphase Model
Interfacial tension can be included in the model using an additional pairwise interaction force $\mathbf{F}^{int}_{ij}$, to the momentum balance (2). This interaction force takes the form tartakovsky2016pairwise
$\mathbf{F}^{int}_{ij}=\mathbf{F}^{int}(\mathbf{r}_{ij})=-\phi(r_{ij})s_{ij}{{\bf e}_{ij}},$ (11)
where $\phi$ and $s_{ij}$ determine the functional form of the potential and the strength, respectively. This force is short-range repulsive and long-range attractive lei2016smoothed . The shape function, $\phi(r_{ij})$, is given by
$\phi(r_{ij})$ $=r_{ij}\left[-Ae^{-{r^{2}_{ij}}/{2\epsilon_{0}^{2}}}+e^{-{r^{2}_{ij}}/{2\epsilon^{2}}}\right],$ (12)
where $A$ represents the magnitude of the ratio between repulsive and attractive forces, and $\epsilon$ and $\epsilon_{0}$ are compact-support dependent functions. For two different phases 1 and 2 the factor $s_{ij}$ is related to the domains $\Omega_{1}$ and $\Omega_{2}$ as
[…]
@@
$s_{11}=s_{22}=10^{k}s_{12}=\frac{1}{1-10^{-k}}d^{-2}_{eq}\frac{\sigma_{0}}{\sqrt{2\pi}(-A\epsilon_{0}^{5}+\epsilon^{5})}$ (14)
Here, $k$ is a constant that must be greater than 1, $\sigma_{0}$ is a macroscopic surface tension and $d_{eq}=1/\Delta x^{D}$ is the equilibrium particle number density.
Following Tartakovsky and Panchenko, we use $\epsilon=h/3.5$, $\epsilon_{0}=\epsilon/2$ and $A=(\epsilon/\epsilon_{0})^{3}$. This choice ensures that the virial pressure ($P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})d_{eq}^{2}s_{11}$) is negative, which is necessary for the stability of the interfaces tartakovsky2016pairwise . However, we must note, that the presence of a negative virial pressure can also lead to numerical instabilities in the system, especially when the system is not in equilibrium. In order to prevent this viral pressure affects the global system pressure, we introduce an equation of state fo the form
@@
where $\rho_{0}$ is the equilibrium density for the system, $c_{s}$ is the speed of sound, and $P_{b}=-P_{v}=-\pi(-A\epsilon_{0}^{4}+\epsilon^{4})d_{eq}^{2}s_{11}$ is a background pressure that ensures $P>0$. This component balance the system and avoid spurious effects of global negative pressure. It is important to note that the presence of this background pressure can also influence the dynamics of the system, especially in flow conditions. In the Results section, we show that the choice of $P_{b}$ does not introduce artifacts into the simulations, under different dynamic conditions.
In our simulations, we adopt the well-known Velvet-velocity algorithm allen1987computer for the temporal evolution of the position and velocity of the particles. We use a time step $\Delta t$ that ensure numerical stabilitybian2012multiscale , satisfying both: $\Delta t\leq 0.125h^{2}\rho/\eta$ (the viscous diffusion criteria morris1997modeling ) and $\Delta t\leq 0.25h/c_{s}$ (the Courant-Friedrichs-Lewy (CFL) condition courant1928 ).
2.3 Relevant time scale
In multiphase thixotropic systems their behavior can be influenced by the interplay of various temporal scales. To streamline the analysis and interpretation of our numerical results, we introduce a set of relevant time scales. The sonic time scale $\tau_{\text{c}}=R/c_{s}$ represents the time it takes for sound waves to propagate across a characteristic length scale $R$. The viscous dissipation time scale $\tau_{\text{vis}}=\rho R^{2}/\eta$, which characterizes the time required for momentum to diffuse across a characteristic length scale $R$ (e.g., droplet radius). The characteristic flow time scale $\tau_{\text{flow}}=R/U$ represents the time it takes for a fluid element to traverse a characteristic length scale $R$ at a velocity $U$. A time scale associated with interfacial tension effects is given by $\tau_{\sigma}=\eta R/\sigma$, which characterizes the time it takes for surface tension to influence the flow over a characteristic length scale $R$. The thermal fluctuations time scale can be estimated from the thermal velocity $v_{th}=\sqrt{k_{B}T/m}$ as $\tau_{\text{t}}=R/v_{th}$, which represents the time it takes for thermal fluctuations to affect the motion of particles over a characteristic length scale $R$. The following hierarchy of time scales holds for realistic multiphase thixotropic systems $\tau_{\text{c}}\ll\tau_{\text{vis}},\tau_{\sigma}\ll\tau_{\text{flow}},\lambda_{\text{thix}},\tau_{\text{t}}$. This hierarchy indicates that sound waves propagate faster than viscous diffusion and interfacial tension, which in turn is faster than convective transport, thermal fluctuations, and microstructural evolution. A summary of the relevant time scales is compiled in Table 1. Our goal is to investigate the interplay between $\tau_{\sigma}$, $\tau_{\text{flow}}$, $\lambda_{\text{thix}}$, and $\tau_{\text{t}}$ time scales in various multiphase thixotropic systems.
From the ratio between these time scales, we can define several dimensionless numbers that characterize the behavior of multiphase thixotropic systems. The Reynolds number is given by $\text{Re}=\tau_{\text{vis}}/\tau_{\text{flow}}=\rho UR/\eta$, which represents the ratio of inertial to viscous forces. The Capillary number is defined as $\text{Ca}=\tau_{\text{flow}}/\tau_{\sigma}=\eta U/\sigma$, which characterizes the relative importance of viscous forces to surface tension forces. The Peclet number is given by $\text{Pe}=\tau_{\text{t}}/\tau_{\text{flow}}=UR/D_{th}$, where $D_{th}$ is the thermal diffusivity, representing the ratio of convective to diffusive transport. Following mujumdar2002transient , larson2019review , rossi2022sph we also defined the so called Thixotropic number as $\text{Th}=(k_{d}/k_{f})(1/\tau_{\text{flow}})$, which evaluates the interplay between the destruction (due to flow) and formation of the microstructure. Where, $\text{Th}>>1$ indicates that microstructural destruction due to the flow is faster than formation, while $\text{Th}<<1$ indicates the opposite. In this work, we focus on conditions where the times scales of formation and destruction of the microstructure overlap $\text{Th}\simeq 1$.
Time Scale Definition Dimensional Analysis
Viscous Diffusion Time Scale $\tau_{\text{vis}}=\frac{\rho R^{2}}{\eta}$ $\frac{[ML^{-3}][L^{2}]}{[ML^{-1}T^{-1}]}=[T]$
Sonic Time Scale $\tau_{\text{c}}=\frac{R}{c_{s}}$ $\frac{[L]}{[LT^{-1}]}=[T]$
Convective Time Scale $\tau_{\text{flow}}=\frac{R}{U}$ $\frac{[L]}{[LT^{-1}]}=[T]$
Interfacial Tension Time Scale $\tau_{\sigma}=\frac{\eta R}{\sigma}$ $\frac{[ML^{-1}T^{-1}][L]}{[MT^{-2}]}=[T]$
Thermal Fluctuation Time Scale $\tau_{\text{t}}=\frac{R}{v_{th}}$ $\frac{[L]}{[LT^{-1}]}=[T]$
Thixotropic Time Scale $\lambda_{\text{thix}}=\frac{1}{k_{f}}$ $k_{f}]=[T^{-1}]\;\Rightarrow\;\lambda_{\text{thix}}=[T]$
Table 1: Relevant time scales in multiphase thixotropic systems with their dimensional analysis, confirming that each has dimensions of time [T].